OpenStudy (sophadof):

Complete each square. A.x^2+10x+16=0 B.x^2+6=-8

jagr2713 (jagr2713):

Hey so what do you think we do first for A

OpenStudy (sophadof):

set up some equation?

OpenStudy (sophadof):

I just don't know what equation :/ @jagr2713

jagr2713 (jagr2713):

I think we are suppose to solve them one by one correct?

OpenStudy (sophadof):

yes! Sorry...I have no idea what these mean.. :( @jagr2713

jagr2713 (jagr2713):

Well what do you think we do first for part A

OpenStudy (sophadof):

Don't we have to divide some variables first? I really don't know how to do this.... @jagr2713

jagr2713 (jagr2713):

So we are factoring. the =0 is something to throw you off but it comes in last so just cover that part out. Do you know how to factor

OpenStudy (sophadof):

yes I think so!

OpenStudy (sophadof):

@jagr2713

jagr2713 (jagr2713):

So what do we do?

jagr2713 (jagr2713):

do you need an explanation on factoring?

OpenStudy (sophadof):

no I will try to figure it out!!!(:

jagr2713 (jagr2713):

k

OpenStudy (sophadof):

so we begin with the regular equation, x^2+10x+16=0 then take the zero off, x^2+10x=16?

OpenStudy (sophadof):

@jagr2713

jagr2713 (jagr2713):

Nope lets leave the zero alone for now x2+10x+16 so when factpr you look at the last number and see if its + or - if its + like it is, find a number that multiples to get the certain number and adds or subtracts up to get the second number.

jagr2713 (jagr2713):

factoring*

OpenStudy (sophadof):

8? @jagr2713

jagr2713 (jagr2713):

and

OpenStudy (sophadof):

2!! @jagr2713

jagr2713 (jagr2713):

Yes so its x^2 we split it into (x+8)(x+2) + means same - means different so so if its for example x^2+6x-5 it is different but still + and minus. It is only + and + when the whole thing is + it is only - and - when its - and + because + means same

jagr2713 (jagr2713):

wow this is a lo of teacher so now we ahev (x+8)(x+2)=0 so now we combine 0 to both sides so we have x+8=0 x+2=0 what do we get?

OpenStudy (sophadof):

I don't wanna waste your time!! I will just try to figure it out myself!!...THank you though, im sorry i aalready wasted some of your time @jagr2713

jagr2713 (jagr2713):

No, i got time and we are almost done

OpenStudy (sophadof):

wait...! I got it! -2,-8 is the answer @jagr2713 haha it wasnt so hard after all

jagr2713 (jagr2713):

Correct how about the second one

OpenStudy (sophadof):

give me a couple of minutes(: @jagr2713

jagr2713 (jagr2713):

k

OpenStudy (sophadof):

(-2,-4) @jagr2713

jagr2713 (jagr2713):

correct

OpenStudy (sophadof):

YAY! thank you so much!! @jagr2713

jagr2713 (jagr2713):

Anytim